40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 7 15:46:42 2013, Jenne, Update, Locking, PR2 flipped 
    Reply  Fri Feb 8 02:12:14 2013, Jenne, Update, Locking, PRMI work 
       Reply  Fri Feb 8 02:38:04 2013, Koji, Update, Locking, PRMI work 
       Reply  Fri Feb 8 11:01:18 2013, Jamie, Update, Locking, PRMI work 
          Reply  Fri Feb 8 12:39:32 2013, yuta, Update, Locking, PRMI work 
          Reply  Fri Feb 8 15:53:48 2013, Jenne, Update, Locking, PRMI work 
             Reply  Fri Feb 8 17:41:34 2013, Jenne, Update, Locking, PRMI work 
                Reply  Fri Feb 8 19:39:02 2013, Koji, Update, Locking, PRMI work 
                Reply  Fri Feb 8 20:05:15 2013, Jenne, Update, Locking, PRMI work 
Message ID: 8042     Entry time: Fri Feb 8 19:39:02 2013     In reply to: 8039
Author: Koji 
Type: Update 
Category: Locking 
Subject: PRMI work 

It seems that the cavity trans looks much better than before. Cool.

At least the optical gain is ~x5 of the previous value. This means what we did was something good.

Looking forward to seeing the further analysis of the caivty...

ELOG V3.1.3-